# Achieving more human brain-like vision via human EEG representational alignment Zitong Lu 1,*, Yile Wang 2, Julie D. Golomb1

### v2

@@ 4.1 Human EEG data for representational alignment
It is worth noting that the training and test sets do not overlap in terms of object categories (concepts), which means that the performance of ReAlnet trained on the training set, when evaluated on the test set, can effectively reveal the model’s generalization capability across different object categories.
@@ 4.2 Human fMRI data for cross-modality testing
To demonstrate that our approach of aligning with human EEG not only enhances the model’s similarity to human EEG but indicates that ReAlnet has effectively learned the human brain’s representational patterns more broadly, we also performed cross-modal testing, testing ReAlnet on data from a different modality (fMRI), from a different set of subjects, viewing a different set of images. The fMRI data originate from [34]. This Shen fMRI dataset recorded human brain fMRI signals from three subjects while they focused on the center of the screen viewing images. We selected the test set from Shen fMRI dataset, which comprises fMRI signals of each subject viewing 50 natural images of different categories from ImageNet, 40 artificial shape images, and 10 alphabetical letter images with each image being viewed 24, 20, and 12 times respectively. We averaged the fMRI signals across the repeated trials to obtain more stable brain activity for each image observation and extracted signals from five regions-of-interest (ROIs) for subsequent comparison of model and human fMRI similarity: V1, V2, V3, V4, and the lateral occipital complex (LOC).
@@ 4.3 Image-to-brain encoding-based alignment pipeline
Basic architecture of ReAlnet and ReAlnet-R: We have chosen the state-of-the-art CORnet-S model [32, 9] as the foundational architecture for ReAlnet, incorporating recurrent connections akin to those in the biological visual system and proven to more closely emulate the brain’s visual processing. Both CORnet and ReAlnet consist of four visual layers (V1, V2, V4, and IT) and a category decoder layer. Layer V1 performs a 7 7 convolution with a stride of 2, followed by a 3 3 max pooling with a stride of 2, and another 3 3 convolution. Layer V2, V4, and IT each perform two 1 1 convolutions, a bottleneck-style 3 3 convolution with a stride of 2, and a 1 1 convolution. Apart from the initial Layer V1, the other three visual layers include recurrent connections, allowing outputs of a certain layer to be passed through the same layer several times (twice in Layer V2 and IT, and four times in Layer V4). We have also chosen another widely used model in image recognition, ResNet18, as the foundational architecture to obtain human EEG-aligned ReAlnet-R, which consists of 18 layers (the last layer is the final decoder to output the predicted category label).
EEG generation module: For ReAlnet, in addition to the original recurrent convolutional neural network structure, we have added an EEG generation module designed to construct an image-to-brain encoding model for generating realistic human EEG signals. Each visual layer is connected to a nonlinear N 128 layer-encoder (Enc-V1, Enc-V2, Enc-V4, and Enc-IT correspond to Layer V1, V2, V4, and IT) that processes through a fully connected network with a ReLU activation. These four layer-encoders are then directly concatenated to form an N 512 Multi-Layer Visual Encoder, which is subsequently connected to an N 340 EEG encoder through a linear layer to generate the predicted EEG signals. Here N is the batch size. For ReAlnet-R which is highly similar to ReAlnet, we extracted the features from Layer 5, Layer 9, Layer 13, and Layer 17 to connect to four nonlinear N 128 layer-encoders through fully connected networks with ReLU activations, and these four layer-encoders are then directly concatenated to form an N 512 Multi-Layer Visual Encoder, which is subsequently connected to an N 340 EEG encoder through a linear layer to generate the predicted EEG signals.
Therefore, we aim for the model to not only perform the object classification task but also to generate human EEG signals which can be highly similar to the real EEG signals when a person views the certain image through the EEG generation module with a series of encoders. During this process of generating brain activity, ReAlnet(-R)’s visual layers are poised to effectively extract features more aligned with neural representations.
@@
Here, represents the -th image, and represents the probability that model predicts the -th image belongs to class out of 1000 categories. However, the correct ImageNet category labels for images in THINGS dataset are not available. Therefore, we adopt the same strategy as in [26], using the labels obtained from the ImageNet pre-trained CORnet without neural alignment as the true labels to stabilize the classification performance of ReAlnet.
is the generation loss, which includes a mean squared error (MSE) loss and a contrastive loss between the generated and real EEG signals. This contrastive loss is calculated based on the dissimilarity (1 minus Spearman correlation coefficient) between generated and real signals, aiming to bring the generated signals from the same image (positive pairs) closer to the corresponding real human EEG signals and make the generated signals from different images (negative pairs) more distinct. is calculated as followed:
@@
Training procedures: Unlike CORnet and ResNet18 which trained on purely image-based ImageNet dataset, ReAlnet and ReAlnet-R additionally trained on individual EEG data. According to ten subjects in THINGS EEG2 dataset, we obtained ten personalized ReAlnets. Each network was trained to minimize the alignment loss including both classification and generation losses with a static loss weight of 100 and a static training rate of 0.00002 for 30 epochs using the Adam optimizer. We used a batch size of 16, meaning the contrastive loss computed dissimilarities of 256 pairs for each gradient step.
Additionally, for ReAlnet, we applied other three different weights ( = 1, 10, or 1000) separately to train the model to further explore the impact of this value on the performance of ReAlnet. We observed that with an increase in , ReAlnets show greater similarity to human EEG and fMRI and more pronounced individual variability within models. However, only ReAlnet with = 100 show significantly higher similarity to human behaviors. Thus, we suggest that = 100 could be the best parameter to conduct the human EEG alignment. Figures S6, S7, S8, S9 and S10 show the performance and similarity results of ReAlnets with different values.
We tested the classification accuracy of ReAlnets on ImageNet at different values (Figure S6). Importantly, to ascertain that the observed decrease in accuracy was not due to the additional generation task compromising classification performance, but rather the absence of correct ImageNet labels for images in THINGS EEG2 dataset, we trained a ReAlnet with = 0. This ReAlnet excluded the EEG signal generation module but underwent fine-tuning with images from THINGS EEG2 dataset. The results indicated that the ReAlnet with = 0 also experienced a similar level of decline.
@@ 4.4 Model-human similarity measurement
To evaluate the similarity between models and human EEG, the shape of each RDM is 200 200, corresponding to 200 images in THINGS EEG2 test set. For EEG RDMs, we applied decoding accuracy between two image conditions as the dissimilarity index to construct EEG RDM for each timepoint and each subject. For model RDMs, we input 200 images into each model and obtained latent features from each visual layer. Then, we constructed each layer’s RDM by calculating the dissimilarity using 1 minus Pearson correlation coefficient between flattened vectors of latent features corresponding to any two images. To compare the representations, we calculated the Spearman correlation coefficient as the similarity index between layer-by-layer model RDMs and timepoint-by-timepoint neural EEG RDMs.
To evaluate the similarity between models and human fMRI, the shape of each RDM (natural images, artificial shape images, or alphabetical letter images) is 50 50, 40 40, or 10 10 in Shen fMRI dataset test set. For fMRI RDMs, we calculated 1 minus Pearson correlation coefficient between voxel-wise activation patterns corresponding to any two images as the dissimilarity index in the RDM for each ROI and each subject. For model RDMs, similar to the EEG comparisons above, we obtained the RDM for each layer from each model. Then, we calculated the Spearman correlation coefficient as the similarity index between layer-by-layer model RDMs and neural fMRI RDMs for different ROIs, assigning the final similarity for a certain brain region as the highest similarity result across model layers due to the lack of a clear correspondence between different model layers and brain regions. All RSA analyses were implemented based on NeuroRA toolbox [49].
@@
Brain-Score is a framework that evaluates how similar artificial neural networks (ANNs) are to the primate visual system [35]. To measure the behavioral similarity between ReAlnet and humans (and monkeys) in visual recognition tasks, we used two behavioral benchmarks from the Brain-Score framework:(https://github.com/brain-score/vision). ”Rajalingham2018public-i2n” assesses the ability of recognizing core objects from visual images, even with various changes in position, size, viewing angle, and background of the objects[7]. ”Geirhos2021-error_consistency” measures the similarity of errors made by ANN and human. Here, metrics of error consistency were adopted in our study, which measure whether there is above-chance overlap in the specific images that humans and models classify incorrectly. [36]. The behavioral Brain-Score is calculated by taking the average of two behavioral benchmarks. We compared the results from ReAlnets and ReAlnet-Rs to the behavioral Brain-Score of CORnet and ResNet respectively, using the same benchmarks. For more detailed information about the behavioral benchmarks used in this study, please refer to the original papers by [35, 7, 36].

### v3

@@ 4.1 Human EEG data for representational alignment
It is worth noting that the training and test sets do not overlap in terms of object categories (concepts), which means that the performance of ReAlnets trained on the training set, when evaluated on the test set, can effectively reveal the model’s generalization capability across different object categories.
@@ 4.2 Human fMRI data for cross-modality testing
To demonstrate that our approach of aligning with human EEG not only enhances the model’s similarity to human EEG but indicates that ReAlnets have effectively learned the human brain’s representational patterns more broadly, we also performed cross-modal testing, testing ReAlnets on data from a different modality (fMRI), from a different set of subjects, viewing a different set of images. The fMRI data originate from [36]. This Shen fMRI dataset recorded human brain fMRI signals from three subjects while they focused on the center of the screen viewing images. We selected the test set from Shen fMRI dataset, which comprises fMRI signals of each subject viewing 50 natural images of different categories from ImageNet, 40 artificial shape images, and 10 alphabetical letter images with each image being viewed 24, 20, and 12 times respectively. We averaged the fMRI signals across the repeated trials to obtain more stable brain activity for each image observation and extracted signals from five regions-of-interest (ROIs) for subsequent comparison of model and human fMRI similarity: V1, V2, V3, V4, and the lateral occipital complex (LOC).
@@ 4.3 Image-to-brain encoding-based alignment pipeline
Basic architecture of ReAlnets and ReAlnet-Rs: We have chosen the state-of-the-art CORnet-S model [34, 9] as the foundational architecture for ReAlnets, incorporating recurrent connections akin to those in the biological visual system and proven to more closely emulate the brain’s visual processing. Both CORnet and ReAlnets consist of four visual layers (V1, V2, V4, and IT) and a category decoder layer. Layer V1 performs a 7 7 convolution with a stride of 2, followed by a 3 3 max pooling with a stride of 2, and another 3 3 convolution. Layer V2, V4, and IT each perform two 1 1 convolutions, a bottleneck-style 3 3 convolution with a stride of 2, and a 1 1 convolution. Apart from the initial Layer V1, the other three visual layers include recurrent connections, allowing outputs of a certain layer to be passed through the same layer several times (twice in Layer V2 and IT, and four times in Layer V4). We have also chosen another widely used model in image recognition, ResNet18, as the foundational architecture to obtain human EEG-aligned ReAlnet-Rs, which consists of 18 layers (the last layer is the final decoder to output the predicted category label).
EEG generation module: For ReAlnets, in addition to the original recurrent convolutional neural network structure, we have added an EEG generation module designed to construct an image-to-brain encoding model for generating realistic human EEG signals. Each visual layer is connected to a nonlinear N 128 layer-encoder (Enc-V1, Enc-V2, Enc-V4, and Enc-IT correspond to Layer V1, V2, V4, and IT) that processes through a fully connected network with a ReLU activation. These four layer-encoders are then directly concatenated to form an N 512 Multi-Layer Visual Encoder, which is subsequently connected to an N 340 EEG encoder through a linear layer to generate the predicted EEG signals. Here N is the batch size. For eAlnet-Rs which is highly similar to ReAlnets, we extracted the features from Layer 5, Layer 9, Layer 13, and Layer 17 to connect to four nonlinear N 128 layer-encoders through fully connected networks with ReLU activations, and these four layer-encoders are then directly concatenated to form an N 512 Multi-Layer Visual Encoder, which is subsequently connected to an N 340 EEG encoder through a linear layer to generate the predicted EEG signals.
Therefore, we aim for the model to not only perform the object classification task but also to generate human EEG signals which can be highly similar to the real EEG signals when a person views the certain image through the EEG generation module with a series of encoders. During this process of generating brain activity, ReAlnet(-R)s’ visual layers are poised to effectively extract features more aligned with neural representations.
@@
Here, represents the -th image, and represents the probability that model predicts the -th image belongs to class out of 1000 categories. However, the correct ImageNet category labels for images in THINGS dataset are not available. Therefore, we adopt the same strategy as in [25], using the labels obtained from the ImageNet pre-trained CORnet without neural alignment as the true labels to stabilize the classification performance of ReAlnets.
is the generation loss, which includes a mean squared error (MSE) loss and a contrastive loss between the generated and real EEG signals. To compute the contrastive loss, we originally aimed to use Spearman correlation to measure the similarity between predicted and target signals. However, as Spearman correlation involves a ranking operation that is non-differentiable, we replaced it with Pearson correlation, which is a differentiable similarity metric. Specifically, the dissimilarity index was calculated as 1 minus Pearson correlation. This substitution allowed us to compute gradients effectively during backpropagation, ensuring the compatibility of the contrastive loss with gradient-based optimization. The contrastive loss aims to bring the generated signals from the same image (positive pairs) closer to the corresponding real human EEG signals and make the generated signals from different images (negative pairs) more distinct. is calculated as followed:
@@
Training procedures: Unlike CORnet and ResNet18 which trained on purely image-based ImageNet dataset, ReAlnets and ReAlnet-Rs additionally trained on individual EEG data. According to ten subjects in THINGS EEG2 dataset, we obtained ten personalized ReAlnets. Each network was trained to minimize the alignment loss including both classification and generation losses with a static loss weight of 100 and a static training rate of 0.00002 for 30 epochs using the Adam optimizer. We used a batch size of 16, meaning the contrastive loss computed dissimilarities of 256 pairs for each gradient step.
Additionally, for ReAlnets, we applied other three different weights ( = 1, 10, or 1000) separately to train the model to further explore the impact of this value on the performance of ReAlnets. We observed that with an increase in , ReAlnets show greater similarity to human EEG and fMRI and more pronounced individual variability within models. However, only ReAlnets with = 100 show significantly higher similarity to human behaviors. Thus, we suggest that = 100 could be the best parameter to conduct the human EEG alignment. Figures S13, S14, S15 and S16 show the performance and similarity results of ReAlnets with different values.
We tested the classification accuracy of ReAlnets on ImageNet at different values (Figure S1A). Importantly, to ascertain that the observed decrease in accuracy was not due to the additional generation task compromising classification performance, but rather the absence of correct ImageNet labels for images in THINGS EEG2 dataset, we trained a ReAlnet with = 0. This ReAlnet excluded the EEG signal generation module but underwent fine-tuning with images from THINGS EEG2 dataset. The results indicated that the ReAlnet with = 0 also experienced a similar level of decline.
Control experiments: To systematically evaluate how different losses and how shuffled EEG data influence our results we conducted four control experiments. The corresponding four control conditions are defined as follows: (1) W/o ContLoss control: We removed the constrastive loss component; (2) W/o MSELoss control: We removed the MSE loss component; (3) Unpaired EEG control: We randomly reassigned EEG signals to images, breaking the image-EEG pairing while preserving the temporal structure of the EEG data; (4) Scrambled EEG control: We temporally permuted the EEG signal within each pair, disrupting the temporal structure of the EEG signal while preserving its overall distribution.
@@ 4.4 Model-human similarity measurement
To evaluate the similarity between models and human EEG, we calculated EEG RDMs using classification-based decoding accuracy as the dissimilarity index. While fMRI RDMs are typically calculated using 1 minus correlation coefficient (below) [61, 62, 63, 2], decoding accuracy is more commonly used for EEG RDMs [64, 65]. Since EEG has a low SNR and includes rapid transient artifacts, Pearson correlations computed over very short time windows yield unstable dissimilarity estimates [66, 67] and may thus fail to reliably detect differences between images. In contrast, decoding accuracy - by training classifiers to focus on task-relevant features - better mitigates noise and highlights representational differences. For each image, we extracted the corresponding EEG response matrix (80 trials 17 channels 20 time-points). For each subject and each time-point, we referred 80 trials per image as training samples and 17 channels as features, resulting in a total of 160 samples for every pair of two images i and j. A linear SVM classifier was employed to classify the EEG responses between two images using a 5-fold cross-validation approach. In each fold the classifier was trained on 4 folds of the data and tested on the remaining fold, ensuring that the decoding accuracy reflected generalizable patterns. The average decoding accuracy across the 5 folds was computed and used as the dissimilarity measure between images i and j. And this process was repeated for all possible pairs of the 200 images, resulting in a 200 × 200 RDM for each subject and each time point. For model RDMs, we input 200 images into each model and obtained latent features from each visual layer. Then, we constructed each layer’s RDM by calculating the dissimilarity using 1 minus Pearson correlation coefficient between flattened vectors of latent features corresponding to any two images. To compare the representations, we calculated the Spearman correlation coefficient as the similarity index between layer-by-layer model RDMs and timepoint-by-timepoint neural EEG RDMs.
To evaluate the similarity between models and human fMRI, we used the Shen fMRI dataset, which includes fMRI activation patterns for different image categories (natural images, artificial shape images, and alphabetical letter images). We calculated separate RDMs for each category, and the RDM dimensions correspond to the number of images in each category: 50 50 for natural images, 40 40 for artificial shape images, and 10 10 for alphabetical letter images. We used the GLM activation values (beta weights) provided in the open dataset for each image and each voxel, and calculated RDMs based on the common correlation-based method [61, 62, 63, 2], as follows. For each subject and ROI, we defined the voxel-wise activation pattern for each image as the vector of activation values (e.g., if a given ROI had 250 voxels, this would be a 1250 vector for each image). Then, for every pair of images, we computed the dissimilarity index as 1 minus Pearson correlation coefficient between the two voxel-wise activation vectors corresponding to the two images. This was repeated for all image pairs, resulting in a symmetrical dissimilarity matrix (the RDM) for each ROI and each subject. For model RDMs, similar to the EEG comparisons above, we obtained the RDM for each layer from each model. Then, we calculated the Spearman correlation coefficient as the similarity index between layer-by-layer model RDMs and neural fMRI RDMs for different ROIs, assigning the final similarity for a certain brain region as the highest similarity result across model layers due to the lack of a clear correspondence between different model layers and brain regions. All RSA analyses were implemented based on NeuroRA toolbox [68].
@@
Brain-Score is a framework that evaluates how similar artificial neural networks (ANNs) are to the primate visual system [37]. To measure the behavioral similarity between ReAlnets and humans (and monkeys) in visual recognition tasks, we used two behavioral benchmarks from the Brain-Score framework:(https://github.com/brain-score/vision). ”Rajalingham2018public-i2n” [7] task uses grayscale images where objects are manipulated by varying position, size, viewing angle, and background, while ”Geirhos2021-error_consistency” [41] task employs out-of-distribution colorful natural images. Both tasks calculate behavioral similarity between the model and human (and primates) observers using the error consistency method, which measure whether there is above-chance overlap in the specific images that humans and models classify incorrectly. The behavioral Brain-Score is calculated by taking the average of two behavioral benchmarks. We compared the results from ReAlnets and ReAlnet-Rs to the behavioral Brain-Score of CORnet and ResNet respectively, using the same benchmarks. For more detailed information about the behavioral benchmarks used in this study, please refer to the original papers by [37, 7, 41].
Individual variability in ReAlnets: To quantify inter-individual representational variability among the individualized ReAlnets, we computed a variability index at each model layer. Specifically, we extracted the RDM from each individualized ReAlnet (one per subject) for a given layer and then calculated all pairwise Spearman correlations between the RDMs of the ten models. The variability index was defined as the average of one minus Spearman correlation coefficient between two RDMs. A higher value indicates greater variability (i.e., less similarity) across models.
Statistical Analysis:
Statistical significance was determined using two-tailed one-sample t-tests between ten ReAlnets and the single CORnet, or paired t-tests between the ten sets of ReAlnets and control models. We considered as the threshold for significance. To ensure that statistical significance reflects meaningful differences, we also report effect sizes in the supplementary materials.
4.5 Model internal representational analysis
This section describes the methodology used for the analysis in Results - Refined object feature representations in ReAlnets section, where we examined which object feature dimensions were more strongly encoded in ReAlnets compared to CORnet. Specifically, we utilized 49 object feature dimensions from the THINGS dataset [42]. Each object concept is represented along these 49 feature dimensions. Our analysis focused on the 200 images in the test set of the THINGS EEG2 dataset, which were not part of the model’s training data. We applied an RDM-based partial Spearman correlation method for the analysis. First, we computed the RDM for the IT layer of each model, which contains higher-level information, and 49 feature RDMs based on the 200 images by calculating the absolute differences in feature encoding strength between pairs of images as dissimilarity measures. Next, we computed the partial correlation between the model RDM and each feature RDM, while controlling for the other 48 feature RDMs. Finally, we calculated the square of the partial correlation coefficient to determine the variance explained by the model for each object feature dimension and got the top-3 improved feature dimensions of ReAlnets.
5 Data availability
The models and the analysis code can be assessed at https://github.com/ZitongLu1996/ReAlnet.